WRITE THE EQUATION in slope intercept form of the line that is perpendicular to y=-5x+2 and passes through 10, 8
Answer by mananth(16946)

1 y = -5 x + 2
Divide by 1
y = -5 x + 2
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = -5

The slope of a line perpendicular to the above line will be the negative reciprocal 1/5
m1*m2=-1
The slope of the required line will be 1/5

m= 1/5 ,point ( 10 , 8 )
Find b by plugging the values of m & the point in
y=mx+b
8 = 2 + b
b= 6
m= 1/5
The required equation isy = 1/ 5 x + 6
m.ananth@hotmail.ca